Thomas’ Calculus Early Transcendentals 12ed

Homework Assignment for Chapter 05

1. Section 5.1: skip reading, just do problem 22.

2. Section 5.2: problem 41.

3. Section 5.3: problems 5, 49, 71, 73 (read Table 5,4 for the Max-Min Inequality). Problem 87 is optional. Do it if time permits. It gives you a feeling (although does not completely answer) why continuous functions are integrable.

4. Section 5.4: odd numbered problems in 1-33. Problems 35, 77, 81, 84, 89.

5. Section 5.5: problems 35, 41, 43, 53, 57, 59, 67, 69, 79.

6. Section 5.6: problems 33, 43, 112, 117.

7. Chap 5, Additional and Advanced Exercises: problems 5, 6, 7, 8.